# Measurement-induced physics and stochastic dynamics in systems of single-body quasiparticle and many-body microscopic cases

### v2

@@ Abstract
In this article, we focus on the measurement-induced physics and stochastic dynamics in the systems of single-body quasiparticle and many-body microscopic cases. In the former case, it shows a smooth evolution from non-Gaussian (Poission or binomial distribution) with local dephasing to Gaussian with global dephasing (quantum to classical) where in late time there is a global symmetry consist of exponentially many pure states due to the suppressed (stochastic) fluctuation and heating; In the latter case, it shows exponentially many distinct (Krylov) sectors due to the fragmentation and dynamical constrain on Hilbert-space.
@@ I Introduction
It has been studied that a pair of coupled identical Sachdev-Ye-Kitaev (SYK) models with complex fermions provide globally chaotic systems which allowing the quantum mutual information to traverse through the wormhole[1], where the robust ergodicity against the localization is supported by the revival oscillations between the two chaotic subsystems[1, 2]. Similarly, a periodic initial state revivals in a many-body system indicate the nonthermal feature in the quench dynamics up to long-time limit (infinite temperature or high energy density)[142, 4, 22], which is distinct to the conventional many-body scars.
The non-Hermitian dynamics may arises from the quantum dissipation or continuous measurement together with postselected trajectories. The non-Hermitian skin effect with finite dissipation (imaginary) gap in complex spectrum shows area-law entanglement scaling. The skin effect is caused by the nonreciprocal dissipation and thus suppressing the entanglement diffusion. Despite sharing the boundary localization, the edge burst[57] additionally require closing dissipative gap[28, 53, 54], origin from a universal bulk-edge scaling relation. Moreover, the multi-partite entanglement can be effectively manipulated through the non-Hermitian (asymmetry) coupling. Note that in this article, for skin effect, we only refer to the nonreciprocal one, i.e., excluding the reciprocal induced by $Z_{2}$-type symmetry[61, 64], and for coherence, we only refer to the one happen in non-Markov case, such that the coherence field (pattern) corresponding to a closed system in a pure state, instead of the global conservation (topological invariance).
@@
In non-Hermitian (open) system, there is unidirectional disentangling due to the coupling to the external disorder. Then there is possibly be infinite number of bound states (in gapped regime) discretically, among which are mutually uncorrelated and the level statistics follows Gaussian ensemble (random matrices). The only way to reveal such external disorder (or non-Hermitian perturbation) is to continuously monitor a individual quantum trajectory, which is a separable pure state (corresponding to a rank-1 subspace) with the continuously measured outcomes be postselected. Here the "individuality" is guaranteed by conditioning the evolution to the absence of quantum jump (sudden change of state) as well as the exceptional point[12, 13, 14, 15]. Such individual trajectory subject to unitary evolution (by certain Hermitian operator $H$) $|\psi(t)\rangle=\frac{e^{-iHt}|\psi_{0}\rangle}{||e^{-iHt}|\psi_{0}\rangle||}$ is conditioned during the continuously measurement and thus support the long-range entanglement, while the repeated local measurement (monitoring) will suppress the short-range entanglement[44, 39] by alternating the random unitaries and introducing the quantum jump processes accompanied by the density fluctuation.
@@ II Effect of measurement in bipartite configuration
While the method of Lindblad master-equation[16] for dissipative Markovian process, whose measurement outcomes are simply averaged out, is inappropriate[13, 15, 14, 17]. Or we can say that the master equation reduced to the form involving a coherent interacting term and a local dissipation term. In this case with the presence of non-Hermitian skin effect, we can suspect the unique external disorder help the system to retain initial information, in opposite with the quasi-particle effect. This corresponds to the densities of the class of many-body microscopic ($mm$) $\rho_{mm;\{i\}}$, which is a Hermitian matrix and characterize the robustness of initial state revival as well as the bulk disentangling liquid which prohibits the diffusion of quantum mutual information (correlation) among the whole system globally. The counterpart of $\rho_{mm;\{i\}}$ is defined as $\rho_{sq;\{i\}}$, which is a non-Hermitian matrix. Here the subscript $sq$ referes to single-body quasiparticle. For $\rho_{sq}$, the non-Hermitian skin effect is absent, and there is delocalization as well as thermalization due to the non-Hermicity induced by $\mathbb{Z}_{2}$-type field, before the thermal equilibrium where the system becomes exactly Hermitian and local with respect to the unique environmental disorder (thermal bath) which should be distingushed from the above-mentioned external disorder for $\rho_{mm}$ Also, for $\rho_{sq}$, there is a diffusion (nonreciprocal dissipation) of quantum mutual information (carried by heavier carrier) among the whole system globally, and thus there will not be disentangling liquid within the bulk but only be a single disentangling boundary. Consequently, such correlation is system-size-dependent, or, extensive. This meets the label of single-body quasi-particle whose lifetime entirely controlled by the unique environmental disorder, and it is impossible for the multi-bound-states to appear like the case of $\rho_{mm}$. Importantly, even the non-Hermitian density $\rho_{sq}$ is diagonalizable due to the existence of non-defective degeneracy (NDD). While for the non-Hermitian operators that contains only the defective degeneracy, it becomes nondiagonalizable at certain momentum termed exceptional point. This means the set of eigenvectors that support non-defective degenerate span a full Hilbert space, in contrast to the defective degenerated one. Importantly, the non-defective degeneracy of $\rho_{H}$ preclude the symmetry-protected many-body localization and guarantees the global symmetry (chaos), thus its complex eigenenergies is different from that of a quasiparticle in a system with topologically enclosed exceptional point, such that its imaginary part (inversed lifetime) is directly determined by the unique environmental disorder from thermal bath (unlike the regular quasiparticles induced from a quantum jump process), and we further found that its real part of eigenenergies correlated (uncorrelated) with the imaginary part outside (inside) the NDD block. While for regular quasiparticles the imaginary part of eigenenergies close and away from the exceptional points are always related to different sources. It is found that the topologically stablity of exceptional points requires two or higher dimensions for a nontrivial topological index and can only be created or annihilated in pairs[82], thus the defectively-degenerated system containing exceptional points requires additional dimension compares to the non-defectively degenerated one. This is in consistent with the NDD eigenvectors constructed in our configuration, where the forward-scattering-like eigenvectors in degenerated region preventing the formation of many-body localization, and the superpositions of NDD eigenvectors form a null-(sub)space whose correlation inside is extensive and thus guaranteeing the global symmetry. Such extensive correlation, or the dependence on the size of global system, is in analogy with the topological invariant with noncontractible path in momentum space[82], unlike the contractible path for a non-Hermitian system with defective degeneracy (exceptional points enclosed by the path).
@@ II.2 Short-range entanglement and nonequilibrium steady state
For $mm$-case the short-range entanglements (with nonlocal measurement) is necessary where there are coherences with weak dissipation (i.e., the non-Markovian dynamics with robust memory effect) and finite-time thermalization (in terms of energy spectrum, this can be evidenced by the gapped structure and variant number of bound states[47, 10]). Thus different to the above cases of density-bath (decoherenced and disentangled), the strong system-bath coupling produce bound states with high fidelity.
Another phenomenon in the limit of closed system is tunneling as a stochastic process where the non-perturbational resonance at some certain local spatial sites is overwhelming compares to others due to the statistical fluctuations of the on-site disorder. This give rise to the non-Hermiticity through the mixing of non-adjacent eigenstates and all-to-all coupling as can be evidenced by the complex eigenenergies[13, 56].
@@ IV.1 $sq$-case
The $sq$-case and $mm$-case reflect the entanglement additivity and symmetry additivity, respectively. As a result, the former has a much larger measurement rate, and exhibits more pronounced randomness.
@@ IV.2 $mm$-case
In the presence of finite dissipation, while the non-Hermiticity refers to the microscopical continuous measurements with the conditioned evolution without quantum jump, there are discontinuous conditioned evolution with quantum jumps for the whole picture of $mm$-case. Those involved quantum jumps can be directly evidenced by the repeatly merged Loschmidt echo, which also refers to the coherent field (local oscillator) in a system like optical cavity. For $sq$-case (without postselection), there is volume-law entanglement without saturating[135], while the $mm$-case always process a saturating (area law) to linearly in time (volume law) transition with the increasing complexity of coherence configuration, i.e., the set of local Hamiltonian (disorder-free). It is such locality and coherence that guarantee the fidelity (as well as the subsystem purity).
@@ IV.3 Realizations under measurement
Thus there could be transitions between nonunitary and unitary dynamics: As mentioned in Sec.IV.1, a nonunitary dynamics with nonuniform continuous measurement can be guaranteed by a relaxation time shorter than that the waiting time between quantum jumps (or projective measurements)[90, 80], and the absence of interaction effect therein allows the measurement-induced phase transition as well as the criticality phenomenon[90, 118]. The high measurement rate (frequency) cause a short Zeno timescale and coarse-grained dynamics due to the incomparable timescales between system and bath [119, 121, 80, 90, 117, 96, 144, 137]. The fast relaxation to equilibrium corresponds to the small dissipation rate $\gamma_{i}$, low occupation fraction, and small noise (two-point) correlation in non-Markovian limit[96] which consistent with the weak noise limit and the Lindblad master equation dominated by the unperturbed (dissipative and uncoherented) part. Fast relaxation gives rise to the dynamical stability due to the suppressed imaginary part of effective non-Hermitian Hamiltonian, as well as the transition to localized phase at steady state in the long-time limit with uniform projective measurement (and thus becomes less sensitive to the measurement unlike the conditioned or postselected dynamics), where the equilibrium state (ensemble average of trajectories[90] or coherent superpositions of symmetry-breaking states[94]), becomes indistinguishable from the time averaged state[80, 93]. Thus, in conclusion, the $sq$-case involving two stages in evolution: In early stage there are weak but frequent measurements (continuously) and strong system-probe coupling[80], dominated by nonunitary (non-Hermitian) stochastic dynamics and algebraic decaying with gapless nature toward the pure state. In late stage where the system approaches the equilibrium steady state, it becomes more significant for the nonextensive character of the localization length and density imbalance of particle distribution due to the vanishingly small dissipation rate and system-probe coupling, in which case the (group) velocity is well-defined according to the long-range power-law hopping of free fermions[56] due to the absence of many-body feature, and meanwhile it is possible to realize unitary and Hermitian dynamics due to the emergent topological invariants[95] (e.g., the nontrivial $\mathbb{Z}$ class[107]) which suppress the short-range entanglement and thus promote the transition from condensated phase (topologically trivial phase) to the topologically ordered pure state (the product state representing an ensemble). In the case of strong Markovian noise, even the shortest-range (nearest neighbor) incoherent hopping in perturbative treatment would leads to diffusion[120]. Note that, different to the "Hermiticity" of $mm$-case where there is power law decay (also including the Gaussian decay of decoherence by the strong system-bath coupling[115]), the "Hermiticity" in late-stage of $sq$ case exhibits exponential decay where the locality is guaranteed by the large dissipative gap under nonunitary (disentangling) measurements[119, 104, 133], and this corresponds to the macroscopic response to the quantum jumps measured in a timescale (waiting time) larger than the fluctuation (scattering induced by imaginary potential[122]) timescale. Equivalently, the fast relaxation to the (pure) dark state in $sq$-case contributes to the global area law phase (e.g., the $Z_{2}$-type localized phase)[119, 15]. The condensed phase here in $mm$-case does not involve the fluctuation and dissipation stochastic contributions from the quantum jumps, and the power-law decay guarantee the (solvable) nonperturbative characteristic[104, 116, 117, 127] (until the sparse measurements and dissipations as well as perturbative contribution from jumps set in). Despite as a hallmark of valid bosonization and condensation and kinetic constraint in Hilbert space[127, 104, 51], the diffusive power-law scaling of long-range hopping and fluctuations is unique to the $sq$-case (attributed to the ergodicity and weak measurements) [56, 126, 39, 96, 128, 129, 130], e.g., the system with anisotropic power-law long-range dipolar interactions[132, 134, 142, 153] which contributes to the polarization and conservation of collective spin and critically slow thermalization (e.g., slow relaxation of magnetization or spin polarization)[142, 153, 145] in contrast to the fast Markovian dephasing.
@@
Distinct from our above statement in many-body perspective that the $mm$-case is dominated by algebraic decay with power-law short-range correlations (due to the skin effect induced, e.g., by the sparse measurements) and the depinning of particle density caused by unitary dynamics (delocalized phase[15]), each single-particle Gaussian state in $mm$-case exhibits exponentially decaying short-range correlation (with area law entanglement) in the case of logarithmically growing (unbound) correlation length[119], e.g., in massive boson mode, but such area law phase induced by exponentially decaying correlations (together with the macroscopic wave function for photons or condensed bosons, i.e., paired fermions with short-range hopping) only takes a little fraction in the $mm$-case due to the isotropic nature and the encoded robust initial state information as long as the system is dominated by the condensation. This is also in agreement with [136], where the classical stochastic dynamics can be overwhelmed from the many-body microscopic details. Despite the skin effect, $mm$-case is dominated by the algebra (power-law) decaying, and indefinitely heating up toward the infinite temperature and maximally mixed steady state, as can be evidenced by the absence of a characteristic time scale in the long-time dynamics and the presence of significant phase fluctuation (suppressed by the measurements and implies no thermalization[134, 135, 136, 137]; see Fig.) and gapless structure of Hermitian measurement operator or Liouville superoperator[133]. Thus, the entanglement entropy loss (heat loss) induced by the sparse quantum jumps is overwhelmed by the entanglement entropy gain due to the non-Hermitian evolution in each sector as well as the non-Gaussian interactions[136, 134].
@@
Thus in perspective of the whole system, the topological order pure state with long-range entanglement and lower measurement rate is favored in $sq$-case where particle-hole symmetry is broken (irreversible) by the long-range hopping, while disordered phase in constricted Hilberts space with short-range entanglement and higher measurement rate is favored in $mm$-case where particle-hole symmetry is preserved (reversible) by the short-range hopping[104, 122, 106, 107]. The topological nontrivial phase in $sq$-case, as stabilized by engineered dissipations, refers to the global symmetries, like the $\mathbb{Z}$ class, to guarantees the bulk-boundary correspondence or the valid description of winding number; While the topological trivial phase refers to the discrete symmetrie, like the particle-hole symmetry. Such disordered-to-ordered transition is also reminiscent of the strong-to-weak spontaneous symmetry breaking (in opposite to the effect of local decoherence from a short-depth local quantum channel[81, 78, 44, 113, 114, 117, 153]), or the reversible-to-irreversible transition, where a description of bosonization is valid in Hermitian limit ($mm$) as evidenced by the linear infinite temperature average (Gaussian correlation) as well as the unsuppressed heating up and noise-induced fluctuations[104, 112, 119, 116, 121], but breaks down in low-density regime with nonunitary dynamics and nonperturbative treatment under continuous post-selected measurements[104, 119]. The Anderson localization is favored under stronger (static) disorder in the late-time stage with saturated localization length. Thus the $sq$-case with Wiener process (Brownian motion)[97, 98, 90] involving both the deterministic (single trajectory level) and Ito stochastic under weak (dephasing) noise. Despite the full dynamics is nonunitary, the generic nonintegrable system
While a unitary dynamics emerges when the uniformly measurement is performed to ensure constant dissipation rate $\gamma_{i}$, and $\sum_{i}L_{i}^{{\dagger}}L_{i}$ is bounded to be a conserved quantity which is inversely proportional to the waiting time between projective measurements[90].
This is consistent with the fact that even in $mm$-case which is dominated by the Hermitian unitary evolution (mixed state with Lindblad master equation), there are repeatly emergence of long-lived pure state due to the coherence field. Such discrete modes and skin effect prohibit the measurement-induced phase transition as well as the diffusion of quantum information in time and space and cause the area-law entanglement. This corresponds to the stronger measurement (denser in frequency) with each eigenvalue has a distinct imaginary part (exhibits nonunitary and non-Hermiticity)[60]. This corresponds to the above effective non-Hermitian Hamiltonian containing a nonuniform measurement (or dissipation) strength $\gamma_{i}$[90]; While for weak measurement with lower frequency, where the eigenvalue of non-Hermitian Hamiltonian have the same imaginary part, there is volume-law entanglement within each sector of $mm$-case (exhibits unitary and Hermiticity). This again verifies the alternative evolutions of unitary and nonunitary ways for point-process dynamics (random projective measurement) of quantum trajectory.

### v4

@@ Abstract
In this article, we focus on measurement-induced physics and stochastic dynamics in the systems of single-body quasiparticle and many-body microscopic cases. In the former case, it shows a smooth evolution from a non-Gaussian (Poission or binomial distribution) with local dephasing to a Gaussian with global dephasing (quantum microscopic to classical macroscopic) where in late time there is a global symmetry that consists of exponentially many pure states due to the suppressed (stochastic) fluctuation and heating; In the latter case, it shows exponentially many distinct (Krylov) sectors due to the fragmentation and dynamical constrain on Hilbert-space. A gapless-to-gapped transition ((sub)volume law to area-law) during evolution is evident for the former case, where, however, the asymptotic approaching area-law scaling (system-size-independence) even appears in early time with the increasing system-size toward infinity. A novel purifying lasting for a short time is found in the early stage of the former case, during which there is long-range power law decay and scrambling before the Lieb-Robinson bound. Meanwhile, in the latter case, local non-Gaussian interactions (with a large power-law exponent) are present, maintaining an invariant gapless feature. Except a comprehensive study for the both cases, our scheme avoid the interaction effect as well as the selections on measured states due to the measurement apparatus, and thus easier evaluatin on the expectations of non-Hermitian observables.
@@ I Introduction
It has been studied that a pair of coupled identical Sachdev-Ye-Kitaev (SYK) models with complex fermions provide globally chaotic systems which allowing the quantum mutual information to traverse through the wormhole[1], where the robust ergodicity against the localization is supported by the revival oscillations between the two chaotic subsystems[1, 2]. Similarly, a periodic initial state revivals in a many-body system indicate the nonthermal feature in the quench dynamics up to long-time limit (infinite temperature or high energy density)[142, 172, 22], which is distinct to the conventional many-body scars. The non-Hermitian dynamics can arises from the quantum dissipation or continuous measurement together with postselected trajectories (exponentially costly[158, 28]). The non-Hermitian skin effect with finite dissipation (imaginary) gap in complex spectrum shows area-law entanglement scaling. The skin effect is caused by the nonreciprocal dissipation and thus suppressing the entanglement diffusion. Despite sharing the boundary localization, the edge burst[57] additionally require closing dissipative gap[28, 53, 54], origin from a universal bulk-edge scaling relation. Moreover, the multi-partite entanglement can be effectively manipulated through the non-Hermitian (asymmetry) coupling. Note that in this article, for skin effect, we only refer to the nonreciprocal one, i.e., excluding the reciprocal induced by $Z_{2}$-type symmetry[61, 64], and for coherence, we only refer to the one happen in non-Markov case, such that the coherence field (pattern) corresponding to a closed system in a pure state, instead of the global conservation (topological invariance). In non-Hermitian (open) system, there is unidirectional disentangling due to the coupling to the external disorder. Then there is possibly be infinite number of bound states (in gapped regime) discretically, among which are mutually uncorrelated and the level statistics follows Gaussian ensemble (random matrices). The only way to reveal such external disorder (or non-Hermitian perturbation) is to continuously monitor a individual quantum trajectory, which is a separable pure state (corresponding to a rank-1 subspace) with the continuously measured outcomes be postselected. Here the "individuality" is guaranteed by conditioning the evolution to the absence of quantum jump (sudden change of state) as well as the exceptional point[12, 13, 14, 15]. Such individual trajectory subject to unitary evolution (by certain Hermitian operator $H$) $|\psi(t)\rangle=\frac{e^{-iHt}|\psi_{0}\rangle}{||e^{-iHt}|\psi_{0}\rangle||}$ is conditioned during the continuously measurement and thus support the long-range entanglement, while the repeated local measurement (monitoring) will suppress the short-range entanglement[44, 39] by alternating the random unitaries and introducing the quantum jump processes accompanied by the density fluctuation.
In this article, instead of exploring the measurement-induced transition, we study the two extreme cases, the limit of weak but continuous measurement and strong but sparse measurement, such that the measurement apparatus plays little role in our scheme. This also provides great convenience in measuring the non-Hermioan observables, where the perturbations-induced imaginary part due to the measurement apparatus can be ignored and thus the preselection on measurement basis is needless. In other word, the uncertainty and randomness are not due to the selections by measurement apparatus, which mean the Green function here does not contains a self-energy term. In many-body microscopic case, the projective measurement suppress the long-range entanglement but does not cause the selections on measured states due to the effect of unique Markovian bath. Another reason is that we choose a time-independent basis $\langle\rho_{0}|\rho_{0}\rangle$. Consequently, the mapping $|\rho_{0}\rangle\rightarrow|\rho^{\prime}_{0}\rangle$ does not cause any new conservations. This indeed reflects the existence of response to the global quench in many-body microscopic case, just like the complete thermal state in late stage of single-body quasiparticle case, i.e., the common characteristic between the Markovian system in single-body quasiparticle case and the unique Markovian bath in many-body microscopic case. These are proved by the expectation (Green function) and variance of non-Hermitia observable[180] as well as the Jensen-Shannon divergence (JSD) in this article.
@@ II Effect of measurement in bipartite configuration
Lindblad master-equation[16] for dissipative Markovian process, consisting of a coherent interacting term and a local dissipation term, and where measurement outcomes are averaged out, is most appropriate[13, 15, 14, 17] in the absence of detection of quantum jumps, in which case the system jumps to the state $|\psi(t+dt)\rangle=L\sqrt{\gamma dt}|\psi(t)\rangle$ with a probability of $\gamma dt\langle\psi|L^{{\dagger}}L|\psi\rangle$ ($L$ is the jump operator and $\gamma$ is the measurement rate). In this case with the presence of non-Hermitian skin effect, we can suspect the unique external disorder help the system to retain initial information, in opposite with the quasi-particle effect. This corresponds to the densities of the class of many-body microscopic ($mm$) $\rho_{mm;\{i\}}$, which is a Hermitian matrix and characterize the robustness of initial state revival as well as the bulk disentangling liquid which prohibits the diffusion of quantum mutual information (correlation) among the whole system globally. The counterpart of $\rho_{mm;\{i\}}$ is defined as $\rho_{sq;\{i\}}$, which is a non-Hermitian matrix. Here the subscript $sq$ referes to single-body quasiparticle. For $\rho_{sq}$, the non-Hermitian skin effect is absent, and there is delocalization as well as thermalization due to the non-Hermicity induced by $\mathbb{Z}_{2}$-type field, before the thermal equilibrium where the system becomes exactly Hermitian and local with respect to the unique environmental disorder (thermal bath) which should be distingushed from the above-mentioned external disorder for $\rho_{mm}$ Also, for $\rho_{sq}$, there is a diffusion (nonreciprocal dissipation) of quantum mutual information (carried by heavier carrier) among the whole system globally, and thus there will not be disentangling liquid within the bulk but only be a single disentangling boundary. Consequently, such correlation is system-size-dependent, or, extensive. This meets the label of single-body quasi-particle whose lifetime entirely controlled by the unique environmental disorder, and it is impossible for the multi-bound-states to appear like the case of $\rho_{mm}$. Importantly, even the non-Hermitian density $\rho_{sq}$ is diagonalizable due to the existence of non-defective degeneracy (NDD). While for the non-Hermitian operators that contains only the defective degeneracy, it becomes nondiagonalizable at certain momentum termed exceptional point. This means the set of eigenvectors that support non-defective degenerate span a full Hilbert space, in contrast to the defective degenerated one. Importantly, the non-defective degeneracy of $\rho_{H}$ preclude the symmetry-protected many-body localization and guarantees the global symmetry (chaos), thus its complex eigenenergies is different from that of a quasiparticle in a system with topologically enclosed exceptional point, such that its imaginary part (inversed lifetime) is directly determined by the unique environmental disorder from thermal bath (unlike the regular quasiparticles induced from a quantum jump process), and we further found that its real part of eigenenergies correlated (uncorrelated) with the imaginary part outside (inside) the NDD block. While for regular quasiparticles the imaginary part of eigenenergies close and away from the exceptional points are always related to different sources. It is found that the topologically stablity of exceptional points requires two or higher dimensions for a nontrivial topological index and can only be created or annihilated in pairs[82], thus the defectively-degenerated system containing exceptional points requires additional dimension compares to the non-defectively degenerated one. This is in consistent with the NDD eigenvectors constructed in our configuration, where the forward-scattering-like[172] eigenvectors in degenerated region preventing the formation of many-body localization, and the superpositions of NDD eigenvectors form a null-(sub)space whose correlation inside is extensive and thus guaranteeing the global symmetry. Such extensive correlation, or the dependence on the size of global system, is in analogy with the topological invariant with noncontractible path in momentum space[82], unlike the contractible path for a non-Hermitian system with defective degeneracy (exceptional points enclosed by the path).
@@ II.2 Short-range entanglement and nonequilibrium steady state
For $mm$-case the short-range entanglements (with nonlocal measurement) is necessary where there are coherences with weak dissipation (i.e., the non-Markovian dynamics with robust memory effect) and finite-time thermalization (in terms of energy spectrum, this can be evidenced by the gapped structure and variant number of bound states[47, 10]). Thus different to the above cases of density-bath (decoherenced and disentangled), the strong system-bath coupling produce bound states with high fidelity. Another phenomenon in the limit of closed system is tunneling as a stochastic process where the non-perturbational resonance at some certain local spatial sites is overwhelming compares to others due to the statistical fluctuations of the on-site disorder. This give rise to the non-Hermiticity through the mixing of non-adjacent eigenstates and all-to-all coupling as can be evidenced by the complex eigenenergies[13, 56, 173].
@@ IV.1 $sq$-case
The $sq$-case and $mm$-case reflect the entanglement (sub)additivity and symmetry additivity, respectively. As a result, the former has a much larger measurement rate, and exhibits more pronounced randomness.
@@ IV.2 $mm$-case
In the presence of finite dissipation, while the non-Hermiticity refers to the microscopical continuous measurements with the conditioned evolution without quantum jump, there are discontinuous conditioned evolution with quantum jumps for the whole picture of $mm$-case. Those involved quantum jumps can be directly evidenced by the repeatly merged Loschmidt echo, which also refers to the coherent field (local oscillator) in a system like optical cavity. $sq$-case always process a linearly-in-time (volume law) to saturating (area law) transition, $mm$-case (without postselection) process a volume-law entanglement without saturating[135] and exhibit persistent fluctuations with the increasing complexity of coherence configuration, i.e., the set of local Hamiltonian (disorder-free). It is such locality and coherence that guarantee the fidelity (as well as the subsystem purity).
@@ IV.3 Realizations under measurement
Thus there could be transitions between nonunitary and unitary dynamics: As mentioned in Sec.IV.1, a nonunitary dynamics with nonuniform continuous measurement can be guaranteed by a relaxation time shorter than that the waiting time between quantum jumps (or projective measurements)[90, 80], and the absence of interaction effect therein allows the measurement-induced phase transition as well as the criticality phenomenon[90, 118]. The high measurement rate (frequency) cause a short Zeno timescale and coarse-grained dynamics due to the incomparable timescales between system and bath [63, 121, 80, 90, 117, 96, 144, 137]. The fast relaxation to equilibrium corresponds to the small dissipation rate $\gamma_{i}$, low occupation fraction, and small noise (two-point) correlation in non-Markovian limit[96] which consistent with the weak noise limit and the Lindblad master equation dominated by the unperturbed (dissipative and uncoherented) part. Fast relaxation gives rise to the dynamical stability due to the suppressed imaginary part of effective non-Hermitian Hamiltonian, as well as the transition to localized phase at steady state in the long-time limit with uniform projective measurement (and thus becomes less sensitive to the measurement unlike the conditioned or postselected dynamics), where the equilibrium state (ensemble average of trajectories[90] or coherent superpositions of symmetry-breaking states[94]), becomes indistinguishable from the time averaged state[80, 93]. Thus, in conclusion, the $sq$-case involving two stages in evolution: In early stage there are weak but frequent measurements (continuously) and strong system-probe coupling[80], dominated by nonunitary (non-Hermitian) stochastic dynamics and algebraic decaying with gapless nature toward the pure state. In late stage where the system approaches the equilibrium steady state, it becomes more significant for the nonextensive character of the localization length and density imbalance of particle distribution due to the vanishingly small dissipation rate and system-probe coupling, in which case the (group) velocity is well-defined according to the long-range power-law hopping of free fermions[56] due to the absence of many-body feature, and meanwhile it is possible to realize unitary and Hermitian dynamics due to the emergent topological invariants[95] (e.g., the nontrivial $\mathbb{Z}$ class[107]) which suppress the short-range entanglement and thus promote the transition from condensated phase (topologically trivial phase) to the topologically ordered pure state (the product state representing an ensemble). In the case of strong Markovian noise, even the shortest-range (nearest neighbor) incoherent hopping in perturbative treatment would leads to diffusion[120]. Note that, different to the "Hermiticity" of $mm$-case where there is power law decay (also including the Gaussian decay of decoherence by the strong system-bath coupling[115]), the "Hermiticity" in late-stage of $sq$ case exhibits exponential decay where the locality is guaranteed by the large dissipative gap under nonunitary (disentangling) measurements[63, 104, 133], and this corresponds to the macroscopic response to the quantum jumps measured in a timescale (waiting time) larger than the fluctuation (scattering induced by imaginary potential[122]) timescale. Equivalently, the fast relaxation to the (pure) dark state in $sq$-case contributes to the global area law phase (e.g., the $Z_{2}$-type localized phase)[63, 15]. The condensed phase here in $mm$-case does not involve the fluctuation and dissipation stochastic contributions from the quantum jumps, and the power-law decay guarantee the (solvable) nonperturbative characteristic[104, 116, 117, 127] (until the sparse measurements and dissipations as well as perturbative contribution from jumps set in). Despite as a hallmark of valid bosonization and condensation and kinetic constraint in Hilbert space[127, 104, 51], the diffusive power-law scaling of long-range hopping and fluctuations is available in $sq$-case (attributed to the ergodicity and weak measurements) [56, 126, 39, 96, 128, 129, 130], e.g., the system with anisotropic power-law long-range dipolar interactions[132, 134, 142, 153] which contributes to the polarization and conservation of collective spin and critically slow thermalization (e.g., slow relaxation of magnetization or spin polarization)[142, 153, 145] in contrast to the fast Markovian dephasing (e.g., the global depolarizing noise channel[174, 175]).
@@
Distinct from our above statement in many-body perspective that the $mm$-case is dominated by algebraic decay with power-law short-range correlations (due to the skin effect induced, e.g., by the sparse measurements) and the depinning of particle density caused by unitary dynamics (delocalized phase[15]), each single-particle Gaussian state in $mm$-case exhibits exponentially decaying short-range correlation (with area law entanglement) in the case of logarithmically growing (unbound) correlation length[63], e.g., in massive boson mode, but such area law phase induced by exponentially decaying correlations (together with the macroscopic wave function for photons or condensed bosons, i.e., paired fermions with short-range hopping) only takes a little fraction in the $mm$-case due to the isotropic nature and the encoded robust initial state information as long as the system is dominated by the condensation. This is also in agreement with [136], where the classical stochastic dynamics can be overwhelmed from the many-body microscopic details. Despite the skin effect, $mm$-case is dominated by the algebra (power-law) decaying, and indefinitely heating up toward the infinite temperature and maximally mixed steady state, as can be evidenced by the absence of a characteristic time scale in the long-time dynamics and the presence of significant phase fluctuation (suppressed by the sparse measurements and implies no thermalization[134, 135, 136, 137]; see Fig.2) and gapless structure of Hermitian measurement operator or Liouville superoperator[133]. Thus, the entanglement entropy loss (heat loss) induced by the sparse quantum jumps is overwhelmed by the entanglement entropy gain due to the non-Hermitian evolution in each sector as well as the non-Gaussian interactions[136, 134].
@@
Thus in perspective of the whole system, the topological order pure state with long-range entanglement and lower measurement rate is favored in $sq$-case where particle-hole symmetry is broken (irreversible) by the long-range hopping, while disordered phase in constricted Hilberts space with short-range entanglement and higher measurement rate is favored in $mm$-case where particle-hole symmetry is preserved (reversible) by the short-range hopping[104, 122, 106, 107]. The topological nontrivial phase in $sq$-case, as stabilized by engineered dissipations, refers to the global symmetries, like the $\mathbb{Z}$ class, to guarantees the bulk-boundary correspondence or the valid description of winding number; While the topological trivial phase refers to the discrete symmetrie, like the particle-hole symmetry. Such disordered-to-ordered transition is also reminiscent of the strong-to-weak spontaneous symmetry breaking (in opposite to the effect of local decoherence from a short-depth local quantum channel[81, 78, 44, 113, 114, 117, 153]), or the reversible-to-irreversible transition, where a description of bosonization is valid in Hermitian limit ($mm$) as evidenced by the linear infinite temperature average (Gaussian correlation) as well as the unsuppressed heating up and noise-induced fluctuations[104, 112, 63, 116, 121], but breaks down in low-density regime with nonunitary dynamics and nonperturbative treatment under continuous post-selected measurements[104, 63]. The Anderson localization is favored under stronger (static) disorder in the late-time stage with saturated localization length. Thus the $sq$-case with Wiener process (Brownian motion)[97, 98, 90] involving both the deterministic (single trajectory level) and Ito stochastic under weak (dephasing) noise. While a unitary dynamics emerges when the uniformly measurement is performed to ensure constant dissipation rate $\gamma_{i}$, and $\sum_{i}L_{i}^{{\dagger}}L_{i}$ is bounded to be a conserved quantity which is inversely proportional to the waiting time between projective measurements[90].